Brand A fertilizer is 32% phosphorus and Brand B is 18% phosphorus. How much of each must be used to produce 56t of a 24% mixture?
Answer by richwmiller(17219)

a+b=56,
32*a+18*b=24*56
a=56-b
32*(56-b)+18*b=1344
1792-32b+18*b=1344
-14*b=-448
b=32
a=56-b
a=24 ton at 32% b=32 ton at 18%
check
32*24+18*32=24*56
768+576=1344
1344=1344
ok